# AI’s Social Forcefield: Reshaping Distributed Cognition in Human–AI Teams

### v2

@@ 4.1. Study Design and Methods
The goal of our first study is to examine how an AI assistant might influence the cognitive alignment in human teams. For this purpose, we recruited 20 teams of three or four participants (69 subjects total) to work together on a collaborative problem-solving task via video conferencing. Our study is designed as a randomized controlled trial with a 2 $\times$ 2 between-subjects design. We manipulated the information provided by the AI (helpful vs. unhelpful) and voice (human-sounding vs. robotic-sounding) of the AI assistant. After completing the task, participants answered a survey about their perceptions of the task, the team, and the AI assistant (see Appendix A1 for details).
4.1.1. Experimental Task
@@
Participants collaborated on the “Cursed Treasure Puzzle,” a pattern-recognition challenge that requires identifying hidden rules and combining them into a final solution. [^1] Task details available at original source at https://jaylorch.net/puzzles/CursedTreasure/ The puzzle is ideally suited to study collaborative problem-solving in human-AI settings: it requires teams to jointly infer hidden rules under uncertainty, coordinate distributed reasoning, manage iterative hypothesis and testing loops which are reflective of many real-world engineering and research tasks, and integrate partial insights across interdependent subproblems into a shared solution. Furthermore, the task requires communication economy to manage information overload and negotiation of shared understanding given its relies on unfamiliar symbols with no standard names (Clark and Wilkes-Gibbs, 1986; Reagans et al., 2023). These properties make it especially suited to studying emergent coordination, communication, and collective sensemaking—all core components of collaboration (additional justifications for these design choices are provided in Appendix A1). Teams had 40 minutes to work on the puzzle and received one AI-provided clue per milestone.
@@
Figure 2. Three chests from the Cursed Treasure puzzle. A treasure chest is cursed with the “octopus curse” if the gems are separated but not cursed if the gems are touching.:
4.1.2. AI Assistant (Experimental Intervention)
@@
The “AI Puzzle Master” was introduced as a voice-only assistant that provided pre-recorded audio messages at fixed intervals (see recordings here). “Puzzle Master” (sometimes “game master” or “MC”) is a widely used in contexts such as logic puzzle hunts, escape rooms, and Dungeons & Dragons–style games for the neutral facilitator who manages the task and provides feedback, and does not inherently imply social authority or evaluative power. In the helpful condition, messages provided relevant hints (e.g., “The octopus rule relates to the separation between the gems”). In the unhelpful condition, the messages gave misleading or irrelevant clues (e.g., it restarted obvious information about the task that was already known from the tasks instructions). The voice of the AI assistant was manipulated between human (recording of a professional female voice actor) and robotic-sounding (text-to-speech). An external panel validated that the robotic voice was perceived as non-human (see Appendix A1 for details). This controlled design allowed us to isolate the causal effects of AI messages on team processes, while keeping the AI’s role predictable and consistent across groups.
4.1.3. Measuring Linguistic Alignment
@@
We analyzed team conversations to study how groups adopted (or resisted) the AI’s language. Using the full transcripts of all team communication (see Appendix A1 for details on recording and pre-processing steps) we developed a set of terms to track salient objects of the task (i.e., five core curses, plus peripheral terms such as “symbols” and “gems”). Each time a participant referred to a tracked object, we recorded whether they used the AI’s term or another term. Lexical alignment was calculated as the proportion of mentions using the AI’s term out of all mentions for that referent. Values range from 0 (never using the AI’s term) to 1 (always using the AI’s term). This measure allowed us to track how alignment with the AI’s language evolved during the puzzle (see Figure 3).
@@
Figure 3. Illustration of lexical alignment on the AI assistant’s language in one example team.:
4.1.4. Survey Measures
After the puzzle, participants completed surveys on three topics:
  1. (1)
    Shared Mental Models: six-item scale on strategy agreement (Johnson et al., 2007; Mohammed et al., 2010).
  2. (2)
    Trust & AI Perceptions: six items on usefulness and reliability of the AI assistant (Pan and Steed, 2016).
  3. (3)
    AI as a Team Member: single item asking whether the AI was perceived as part of, or leading, the team.
4.1.5. Analysis Method
@@
To estimate the causal effect of AI interventions on language use, we applied difference-in-difference regression models (Angrist and Pischke, 2009). The method compares changes in term usage for objects mentioned by the AI against control objects not mentioned by the AI in the same segment. The causal effect is identified by the inclusion of fixed effects for object, team, and time segment were included, with clustered standard errors at the team level (see full model details and robustness checks in Appendix A1). The analysis captures the causal effect of AI induced changes in language accounting for baseline likelihood of language changes, shifts in topic salience over time, and teams’ overall volume of communication.
4.1.6. Participants
@@ 4.2. Results
This section reports the main findings of Study 1: how the AI assistant influenced the collective attention of teams, both in terms of what they focused on and how they talked about it. We present high-level results here and refer readers to Appendix A1 for detailed case studies, additional survey analyses, and evidence on underlying social and cognitive mechanisms.
4.2.1. Linguistic Alignment
The AI influenced how teams communicated about the objects of discussion, as reflected in the specific word choices they made. Teams generally adopted the AI’s terms after its interventions (Table 2, Model 1: $\beta=0.06$, $p<0.01$), but this effect varied by condition. Alignment was strongest in the Human-Helpful condition and weakest in the Robotic-Unhelpful condition (Table 2).
Dependent Variable: Entrainment on AI Language
@@ 4.2.1. Linguistic Alignment
Why were effects stronger for Human-Helpful? Our survey responses help explain this pattern. Participants rated helpful AIs as more trustworthy and intelligent ($p<0.001$), and they were more likely to see a human-sounding AI as part of their team. Language data further reinforced this: use of the pronoun “we” was highest in the Human-Helpful condition, suggesting the AI was integrated into collective identity and given higher status. In these cases, alignment seems to reflect not only informational influence but also a social process of treating the AI like a valued teammate.
Why the “flipped” effect for peripheral terms? An unexpected effect emerged for peripheral referents, which we define as puzzle elements that were part of the discussion but not essential for finding the solution (Table 2, Model 5). These elements include descriptive or secondary labels (e.g., “symbols,” “gems”) that teams could talk about without directly affecting their progress on the main milestones. In unhelpful conditions, teams resisted the AI’s terminology for core concepts but showed a higher likelihood of adopting them for peripheral ones. This likely reflects a low-stakes channel of influence: adopting the AI’s language for non-essential elements carried little cost and reduced cognitive effort, consistent with prior findings on linguistic alignment in low-stakes contexts (Vasconcelos et al., 2023). In contrast, for core concepts, excluding the AI’s language may have helped teams reinforce their own shared understanding and social cohesion.
4.2.2. Shared Attention & Alignment of Mental Models
@@
The AI assistant had a consistent impact on what teams discussed, shaping their collective focus of attention. Teams were significantly more likely to talk about aspects of the task immediately after the AI introduced it (Figure 4). Using the same difference-in-difference analysis approach confirmed this effect ($\beta=0.35$; $p<0.001$), showing that the AI reliably seeded new discussion topics, regardless of whether its clues were helpful or unhelpful, or whether its voice was human or robotic.
@@
Figure 4. AI affects the shared focus of attention. Teams talk more about the object mentioned by the AI in the segment immediately after its interjection than at other times. Counts include all mentions of tracked objects, regardless of which terms are used.:
@@
The AI also had significant effect on teams shared mental models (one-way ANOVA $F(3,65)=5.5$; $p=0.041$). We find a consistent level of alignment in shared mental models in both helpful conditions, while the unhelpful conditions exhibit diverging patterns (Figure A2). Mental models are less well aligned (i.e., have higher standard deviation) in the Human-Unhelpful condition, while we find the highest level of alignment in the Robotic-Unhelpful condition. The combination of evidence from behavioral measures of group interaction (lexical alignment) and participants’ self-reflection of their group’s process (shared mental models) offers particularly compelling support for the AI assistant’s influence on cognitive alignment (DeChurch and Mesmer-Magnus, 2010; Lix et al., 2022).
4.2.3. Group Identity and Social Cohesion
@@
Finally, we explore explore the AI’s effect on group identity and social cohesion through the use of first- and third-person pronouns (Figure 5; see Appendix A1 for additional method details). Teams in the Human-Helpful and Robotic-Helpful conditions showed a moderate level of pronoun use (a mean of 281 and 269 pronouns per team, respectively). Teams in the Robotic-Unhelpful treatment used considerably more pronouns (mean of 297 per team) indicating much stronger group identity social cohesion, while teams in the Human-Unhelpful treatment used significantly fewer (mean of 223 per team), suggesting little cohesion. In particular “we” is the most frequently used pronoun in the Human-Helpful condition but is used rarely in the Human-Unhelpful condition. This pattern suggests that the presence of an unhelpful, yet human-sounding AI assistant disrupted the natural cognitive alignment in teams and undermined their social cohesion. In the Robotic-Unhelpful condition, the unnatural voice of the AI makes it easier for the human participants to exclude it from their in-group thus strengthening their social cohesion. The failure to exclude the disruptive AI agent also explains the poor alignment of mental models in the Human-Unhelpful condition.
4.2.4. Collective Cognition as Multi-Level Alignment
Taken together, these results show that AI assistants do shape collective cognition not through a single mechanism, but through multi-level alignment. AI induces linguistic alignment of word choices, which further scaffolds shared attentional focus and alignment of mental models, and shifted perceptions of group membership. From a dynamical systems perspective, this fits with theories of human interaction as self-organizing synergies across multiple alignment channels, including joint attention, linguistic entrainment, social identity, and shared intentionality (Hultman et al., 2022; Gallotti et al., 2017). It is important to realize that even a scripted, non-interactive AI can perturb this alignment system, redirecting focus, reshaping language, and altering the social boundaries of teamwork.
@@ 4.2.4. Collective Cognition as Multi-Level Alignment
5. Study 2 - Customer Service Encounters
We complement our first study with a second preregistered [^2] https://osf.io/quj2d/?view_only=9299fa86594b487c9c356e1a32dbe6f7 and IRB approved study. This study uses a minimal design with a state-of-the-art interactive chatbot. It directly tests whether subtle variations in an AI assistant’s framing and language can leave a cognitive and linguistic “fingerprint” that persists when people later communicate with other humans—even when the AI is no longer present. The study design randomly manipulates the AI’s system prompt and measures how much subjects spontaneously reuse AI words in a human-led debriefing. This experimental manipulation of linguistic exposure thus isolates the causal pathway through which AI priming reshapes human language and thought.
Study Design and Task
To achieve high ecological validity, we designed a study around a realistic task—AI assisting customer support agents who perform text-based chat support to resolve customer complaints and technical problems. AI is already widely used for such tasks (BCG, 2024). One prominent field study in a Fortune 500 company showed that AI is highly effective supporting customer service agents, reporting an 14% average increase in productivity (Brynjolfsson et al., 2025). Our study had two phases. First, users interacted with a state-of-the-art chatbot (ChatGPT 4o) (Hurst et al., 2024) to write answers to five customer service complaints across different domains (including technical trouble shooting, online shopping assistance, and reservation booking; one example service complaint was “Every time I try to watch a movie, it buffers endlessly”). The AI was a fully functional ChatGPT 4o assistant. Subjects could interact with it like with any chatbot system, write any prompt they liked, for as many rounds as they wanted. Subjects interacted with the AI on a study-provided laptop computer. Users were randomly assigned to one of two AI conditions. Through the system prompt we instructed ChatGPT to either suggest customer service responses that emphasized warmth and empathy or formal adherence to company guidelines (see full system prompts in Appendix A2). The system prompt manipulation subtly changed which words subjects were exposed to by the AI. We term these exposure words and they provide the key experimental manipulation and basis for our analysis. We recorded all user-written prompts and ChatGPT responses (i.e., exposure words), as well as the ultimate response to the customer complaint through a custom user interface implemented via Qualtrics. Each conversation started with ChatGPT offering a suggested draft response to the customer complaint. Subjects wrote on average one additional prompt per customer complaint to refine the initial draft response generated by the chatbot ($SD=1$). The typical interaction was two prompts long. We find no significant different in engagement between the two treatment conditions (two sample t-test, $t=0.843$, $df=92.988$, $p=0.401$).
After the interaction with the AI, subjects participated in an in-person, face-to-face debriefing interview with an experimenter that lasted about five minutes (SD: 108 seconds). The debriefing interview questions were exclusively focused on the tasks—not the AI. To capture not only how the AI shapes how people talk, but also how they imagine acting in the future—their cognitive-behavioral approach to the task—questions were specifically designed to elicit subject’s process, goal, or approach to the task. For example, one question from the interview guide was “If you were working on a real customer service team, what kind of approach would you suggest the team agree on for answering complaints?” A common follow-up question explicitly elicited subject’s approach to the task and behavioral intention “What’s the first thing you usually focus on [when answering customer service complaints]?” In this way, our measure of overlap captures not only linguistic entrainment but also what aspects of the task are salient and meaningful to people, thus capturing important aspects of cognitive alignment through which groups sustain a common frame of reference for joint action (Shteynberg, 2015; Clark, 1996; Clark and Brennan, 1991; Tomasello, 2009; Hutchins, 1995).
Furthermore, interviews were explicitly framed as team-interaction questions. That is, the conversation was focus on collective group-level constructs, framed around how teams that the subject is part of: the collective unit of the customer service team (see quote above) (Mathieu et al., 2022). This form of “we” narratives are not only expressions shared language, mutual belief, and group identity (Gallagher and Tollefsen, 2019), they help construct and maintain social relations (Clark, 1996). Saying “we” instead of “I” positions the speaker as part of a collective agent and enacts a shared intentional frame. This framing was highly effective and 35% of utterances included the words “we” or “team” (example response from one subject: “[customers] should be let know that we are working on the problem and we resolve it as soon as possible and the team should be assigned as soon as possible to work on it, yeah”—emphasis ours). The conversation thus also reflects important social cohesion and team identity.
Experimenters were blind to the treatment condition and exposure words that subjects experienced during their task work. They were instructed to maintain a neutral, non-judgmental tone, mention the AI’s language explicitly, or that the goal of the study, and instead focus on elicit detailed verbal descriptions of the task the participant worked on and how they thought about it. We recorded the debriefing interview with wireless lapel microphones via smartphones or laptops and used Amazon Web Services Transcribe to transcribe interviews with automated speaker segmentation. We term the words spoken by subjects during the debrief interview spoken words. Overlap between exposure words and spoken words capture the extent to which the influence from the AI assistant spills over to natural human-human conversations and persists even when AI is no longer an active participant of the interaction.
This is a particularly strong and realistic research design. It mirrors a workplace setting in which individuals interact with AI alone during solo work yet they are embedded in a broader human-human team—a social context in which they develop and discuss shared norms of their task, instruct junior team members how to perform the task, and share behavioral strategies they find effective when completing their task. All of these social interactions can be influenced by individual’s experiences and exposure to AI. This tightly controlled design with a realistic, fully functional chatbot, complements the design of Study 1 in which team language was jointly negotiated, measured shared term use across participants, and included measures shared model alignment. How do we analyze this data?
Permutation Test.
We compared the overlap between exposure words subjects saw during their chatbot interaction and their spoken words they later used. Building on the logic of our randomized controlled trial, we contrast the true overlap with an equally likely counterfactual overlap if subjects had been in the opposite treatment condition and had experienced different AI responses. Specifically, for each utterance, we constructed a matched counterfactual: the same utterance, but with exposure words drawn from a randomly sampled participant in the opposite treatment condition. The analysis then compares observed vs. counterfactual counts of overlapping words for the same utterance. If AI exposure has no effect on subjects spoken words, then treatment assignments are exchangeable, and the observed overlap should look identical then the permuted null distribution of overlaps. Formally, we test the null hypothesis: AI exposure has no effect on spoken words. This counterfactual construction provides a rigorous within-utterance (some utterances are longer than others), within-participant (some individuals speak a lot; or are exposed to more words by engaging in more AI turns) control that makes the test more sensitive and interpretable than treating the raw counts in isolation. This analysis isolates the unique treatment signal (observed-counterfactual differences), not just reuse of AI language overall (Table A1 shows example exposure words across treatments). We perform $B=2,000$ permutation shuffles: each iterations draws one counterfactual set of exposure words for each user. We repeat the analysis using both raw overlap counts, and after removing stopwords and stemming and find similar results (we report the stemmed results).
Participants
We recruited 97 subjects on campus at a major private US university from the East Coast. Subjects were compensated with \$10 in Amazon gift card. Average age was 24 [18-34], self-reported gender was 46% female, 78% were students, 13% employed, 7% unemployed, 59% held a bachelor’s degree, 30% a master’s, 9% reported some college education.
Results.
We find AI exposure significantly affects language use. In expectation, per-utterance overlap should be 2.79 words yet observed overlap is 3.00 words ($p=0.0185$, Figure A3). As robustness test, we repeat the analysis but remove utterances in which the subject explicitly referred to the AI. This analysis captures automatic, naturalistic adoption of AI language—AI words slipping into natural speech—beyond direct rephrasing of AI suggested terms (e.g., we remove utterances like “I feel like the AI was, um, […] I want you to show empathy”). Results are consistent ($p=0.021$, Figure A3). As an additional robustness test, we replace the raw counts of overlap as test statistic with the coefficient from a Poisson which allows us to use additional controls and user-level fixed effects (see Eq. 2). This modeling approach controls for opportunity (utterance length), eliminates all person-level confounds such as baseline tendency to reuse any words (i.e., anchoring bias), engagement level, attention paid to the AI, and controls for utterance-specific factors such as topic, emotion, length, difficulty, and timing. We again find significant increase in AI used words ($\beta=0.151$, 95% CI [0.013 - 0.308], $p=0.025$ based on within–between variance decomposition standard error). That is, exposure to AI words increases the likelihood of using those words in subsequent speech by 16% compared to counterfactual words. For a typical 20-word utterance, this corresponds to roughly one additional word from the unique set of AI exposure words from their assigned treatment condition (e.g., the formal chatbot) that are not present in the unique set of words of the other condition (e.g., the empathetic chatbot)—a subtle but systematic imprint on treatment condition specific AI language. The Appendix reports additional analyses mentioned in the study’s pre-registration which provide additional suggestive evidence of AI’s impact on user language. Overall, the study demonstrates that AI systems do not merely assist individuals in the moment but actively and often unconsciously reconfigure how people express themselves and coordinate socially afterward—validating this paper’s broader claim that AI can subtly but powerfully reshape the social fabric of teamwork.

### v3

@@ 5.1. Study Design and Methods
Experimental Task
@@
Participants collaborated for 40 minutes on the a complex, multi-stage pattern-recognition task that requires identifying hidden rules and combining them into a final solution (Fig. 2). [^3] Task details available at original source at https://jaylorch.net/puzzles/CursedTreasure/ The task is ideally suited to study collaborative problem-solving in human-AI settings: it requires teams to jointly infer hidden rules under uncertainty, coordinate distributed reasoning, manage iterative hypothesis testing loops, and integrate partial insights across interdependent subproblems into a shared solution—all of which are reflective of many real-world engineering and research tasks. Furthermore, the task requires communication economy to manage information overload and negotiation of shared understanding given it relies on unfamiliar symbols with no standard names (Clark and Wilkes-Gibbs, 1986; Reagans et al., 2023). These properties make it especially suited to studying emergent coordination, communication, and collective sensemaking—all core components of collaboration (additional justifications for these design choices are provided in Appendix A2).
@@
Figure 2. Three chests from the Cursed Treasure puzzle. A treasure chest is cursed with the “octopus curse” if the gems are separated but not cursed if the gems are touching.
A sample of three chests in the Cursed Treasure puzzle. The octopus curse relates to the separation of gems.
AI Assistant
@@
Teams were assigned an AI-based “Puzzle Master” that was introduced as a voice-only assistant that provided pre-recorded audio messages at fixed intervals (see recordings here). “Puzzle Master” (sometimes “game master” or “MC”) is a widely used term in contexts such as logic puzzle hunts, escape rooms, and Dungeons & Dragons–style games for the neutral facilitator who manages the task and provides feedback, and does not inherently imply social authority or evaluative power. Indeed, the majority of participants did not attribute high capability or authority to the AI assistant (see results of survey-based measures discussed below). In the helpful condition, messages provided relevant hints (e.g., “the octopus rule relates to the separation between the gems”). In the unhelpful condition, the messages gave misleading or irrelevant clues (such as restating obvious information about the task that was already known from the tasks instructions). The voice of the AI assistant was manipulated to be human-sounding (recording of a professional female voice actor) and robotic-sounding (text-to-speech). An external panel validated that the robotic voice was perceived as non-human (see Appendix A2 for details). This controlled design allowed us to isolate the causal effects of AI messages on team processes, while keeping the AI’s role predictable and consistent across groups.
Measuring AI Language Spillover
@@
We analyzed team conversations to study how groups adopted (or resisted) the AI’s language. Using the full transcripts of all team communication (see Appendix A2 for details on recording and pre-processing steps) we developed a set of terms to track salient objects of the task (i.e., five core milestones (curses), plus peripheral terms such as “symbols” and “gems”). Each time a participant referred to a tracked object, we recorded whether they used the AI’s term or a different term. Linguistic spillover from the AI to the human team members was calculated as the proportion of mentions using the AI’s term out of all mentions for that referent. Values range from 0 (never using the AI’s term) to 1 (always using the AI’s term). This measure allowed us to track how aligned the team became with the AI’s language over the course of the task (see Figure 3).
@@
Figure 3. Illustration of lexical spillover from the AI assistant to human members in one example team.
Lexical alignment patterns in an example team.
Measuring AI Influence on Shared Mental Models
After completing the task, participants answered a survey about their perceptions of the task, the team, and the AI assistant (see Appendix A2 for details). Specifically, we measured the teams cognitive alignment via the six-item Shared Mental Models scale on strategy agreement (Johnson et al., 2007; Mohammed et al., 2010). We also measured participants Trust & AI Perceptions via six items on usefulness and reliability of the AI assistant (Pan and Steed, 2016) and perception of AI as a Team Member (single item asking whether the AI was perceived as part of, or leading, the team).
Analysis Method
@@
To estimate the causal spillover of AI introduced language to communication within the team, we applied difference-in-difference regression models (Angrist and Pischke, 2009). This approach isolates the causal spillover of AI while controlling for baseline language changes we would expect from shared task vocabulary and common ground alone. The method compares changes in term usage for objects mentioned by the AI against control objects not mentioned by the AI in the same segment. The causal effect is identified by the inclusion of fixed effects for object, team, and time segment were included, with clustered standard errors at the team level (see full model details and robustness checks in Appendix A2). Intuitively, if a team already uses the same term as used by the AI (e.g., “gems”) before the AI mentions it, there is no causal spillover from the AI. If, on the other hand, a team happens to use a different term (e.g., “diamonds”) but then switches to using the AI term (i.e., switches to calling it “gems”) in the segment immediately following the AI intervention—and not during other segments, and not changing their use of other tracked terms that were not used during this segment—the spillover can causally be attributed to the AI above and beyond baseline linguistic alignment of other terms and during other periods that naturally follow from the shared task and common ground alone. In summary, the analysis captures the causal spillover of AI induced language accounting for baseline likelihood of language changes, shifts in topic salience over time, and teams’ overall volume of communication in naturally occurring spontaneous speech.
Participants
@@ 5.2. Results
We first complement results from the first study with additional evidence of causal AI language spillover in a multi-person team setting. Then we present results on how AI affects collective attention and shared mental models, and finally how it affects group identity and social cohesion. We focus on presenting high-level results here and refer readers to Appendix A2 for detailed case studies and additional results.
Linguistic Spillover
We find significant spillover of specific AI language on team communication. Teams showed a high likelihood to adopt AI language for directly task relevant terms overall (Table 2, Model 1: $\beta=0.06$, $p<0.01$), with subtle variation across conditions. Spillover was strongest in the Human-Helpful condition (Model 4: $\beta=0.15$, $p<0.001$) and weaker when the AI used a robotic voice (Model 2: $\beta=-0.11$, $p<0.05$) or was unhelpful (Model 3: $\beta=-0.08$, $ns$). In unhelpful conditions, teams were unaffected by the AI’s terminology for core concepts (Model 4: $\beta=-0.16$, $p<0.05$) but showed strong spillover for peripheral terms (Model 5: $\beta=0.27$, $p<0.01$).
Dependent Variable: Spillover of AI Language
@@
AI Influence on Collective Attention
@@
The AI assistant had a consistent impact on the collective focus of attention of teams. Teams were significantly more likely to talk about those aspects of the task that were just mentioned by the AI, above and beyond baseline likelihood to talk about those aspects during other segments in which the AI did not mention them (Figure 4). Using the same difference-in-difference analysis approach confirmed this effect ($\beta=0.35$; $p<0.001$), showing that the AI reliably shifted teams’ collective attention.
@@
Figure 4. AI affects the shared focus of attention. Teams talk more about the object mentioned by the AI in the segment immediately after its interjection than at other times. Counts include all mentions of tracked objects, regardless of which terms are used.
What are teams talking about? Teams talk more about the object mentioned by the AI in the segment immediately after its interjection than at other times.
AI Influence on Shared Mental Models
@@
The AI also had significant effect on teams shared mental models (one-way ANOVA $F(3,65)=5.5$; $p=0.041$). We find a consistent level of alignment in shared mental models in both helpful conditions, while the unhelpful conditions exhibit diverging patterns (Figure A3). The presence of the AI assistant disrupted the natural alignment of shared mental models in the Human-Unhelpful condition, while it facilitated higher alignment in the Robotic-Unhelpful condition.
AI Influence on Group Identity and Social Cohesion
@@
Finally, we explore the AI’s effect on group identity and social cohesion through the use of first- and third-person pronouns (Figure 5; see Appendix A2 for additional method details). A striking finding is that AI’s effects on social cohesion did not simply mirror its effects on other alignment channels: while the Robotic-Unhelpful condition showed the lowest linguistic spillover it showed the strongest social cohesion. This presents an “us-vs-them” dynamic in which the presence of an out-group agent strengthened in-group bonds (consistent with Social Identity Theory (tajfel1979integrative) and work on compensatory solidarity under threat (Echterhoff et al., 2009). Furthermore, the Human-Unhelpful shows disrupted social cohesion (in line with the poor alignment of shared mental models). In particular “we” is the most frequently used pronoun in the Human-Helpful condition but is used rarely in the Human-Unhelpful condition. That is, when the unhelpful AI is robotic sounding, the unnatural voice of the AI makes it easier for the human participants to exclude it from their in-group thus strengthening their social cohesion, while in the human-sounding condition, exclusion of the out-group agent is difficult, leaving the team with low social cohesion. This suggests that AI can both undermine alignment and strengthen it and that the socio-cognitive alignment system does not respond uniformly to the introduction of AI; effects on one channel (e.g., linguistic disruption) may co-occur with countervailing effects on another (e.g., affective strengthening).
@@
Conscious vs. Implicit Influence
Critics may say the effects represent only baseline alignment, or happen consciously because the AI is perceived as an influential, high status actor. In this section, we address these counter arguments. First, our difference-in-difference analyses carefully contrasts affected from unaffected terms and affected from unaffected temporal segments, while controlling for teams’ tendency to adopt any language (team-level fixed effects) and topic focus (term-fixed effects). It thus isolates the causal influence of AI interjections above and beyond naturally occurring shifts in topic focus and language co-construction in teams. Second, our finding contrasting spillover on core from peripheral terms suggests implicit influence: while teams are able to consciously resist AI influence on aspects central to the task when the AI is unhelpful, they are still susceptible to AI influence in peripheral areas. Third, the AI influence occurred despite the fact that teams generally did not trust the AI (especially in the unhelpful condition; see Section A2.4.4) and did not consider it a member of their teams (77% of participants responded negatively; see Section A2.4.4). That is, effects are not dependent on (high) trust in the AI or the perceived inclusion of the AI as a team member. Fourth, the effects do not appear to be contingent on high-fidelity, interactive AI (as in Study 1); they arise even with AI of limited capability and low social presence (the scripted, non-interactive AI in Study 2). Taken together, our findings suggest that AI influence in implicit rather than conscious, and happens independent of perceived trustworthiness, social presence, or capability of the AI. These influences can be both positive (as in the case of strong social cohesion when the Robotic-Unhelpful AI is easy to exclude from the team) and negative (as in the case of disrupted alignment of mental models and social cohesion when the Human-Unhelpful AI is difficult to exclude from the team).